# Quadratic exchange equations for Coxeter matroids

### v1

@@ 1 Introduction
Many have studied (see Section 1.1) how well this happy state of affairs replicates for other quotients $\mathbb{G}/\mathbb{P}$ of a reductive group by a parabolic subgroup. In these studies, the first choice to be made is an embedding of $\mathbb{G}/\mathbb{P}$ in a projective space with a prescribed coordinate basis, which is required to define the tropicalisation. Our objective in this paper is to do so uniformly in $\mathbb{G}$. Representation theory gives a uniform procedure to define a projective embedding of $\mathbb{G}/\mathbb{P}$ in a projectivised $\mathbb{G}$-representation $\operatorname{proj}(V_{\lambda})$, in which it is cut out by quadrics (Theorem 4.2). As for the coordinate basis, for Grassmannians we can concisely express the choice as the basis of homogeneous linear forms in the weight grading for $\mathbb{G}$; since the weight spaces are one-dimensional, this uniquely specifies our basis up to scalars. The same approach works for $\mathbb{G}/\mathbb{P}$ when $V_{\lambda}$ is minuscule (we also call $\mathbb{P}$ minuscule in this case), i.e. all the weights of $V_{\lambda}$ are in a single Weyl group orbit. This implies that all weight spaces are one-dimensional. So it is the minuscule case we study here. For convenience in the representation theory, our paper will set $\mathbb{K}=\mathbb{C}$ (although see Remark 4.1).
On the other side of the connection, Coxeter matroids [BorovikGelfandWhite2003] are to $\mathbb{G}/\mathbb{P}$ as matroids are to the Grassmannian. A Coxeter matroid is any subpolytope of the orbit polytope associated to $\mathbb{P}$ of the Weyl group $W(\mathbb{G})$ whose edges are parallel to roots of $\mathbb{G}$ (see Section 2.3 for a careful definition). With this, we have every piece but one of our main theorem:
@@
The one word yet to be defined is “strong”. We say that a Coxeter matroid $P$ is strong if for any two vertices $v,w$ of $P$, either $\operatorname{conv}\{v,w\}$ is parallel to a root or it is the diagonal of a parallelogram all four of whose vertices are vertices of $P$ and with an edge parallel to a root.s (Such Coxeter matroids are also said to satisfy the strong exchange property. For more details see Section 2.3 below.) The condition of being strong will emerge from the structure of tropicalised quadrics homogeneous in the weight grading. In particular, if $P$ is realizable in the sense that there is a point $x\in\mathbb{G}/\mathbb{P}$ so that a vertex $v$ of the orbit polytope is present in $P$ if and only if the weight-$v$ coordinate of $x$ is nonzero, then $P$ satisfies every tropical equation of $\mathbb{G}/\mathbb{P}$, so $P$ is strong. All (ordinary) matroids are strong, but not, for example, all $\Delta$-matroids (Section 2.5). For Coxeter matroids that are not strong, we are unable to suggest any characterisation by equations of tropical hypersurfaces, whatever their degree.
Borovik, Gelfand and White recognised the importance of minuscule representations in Open Problem 6.16.1 of [BorovikGelfandWhite2003]. Our theorem can be read as morally confirming the “if” direction of their expectation in that open problem, strengthened by the connection to equations of $\mathbb{G}/\mathbb{P}$.
@@ Acknowledgements
The authors thank Chris Eur for his part in the initial impetus for this project and for pointing us to [Kumar2002], and Alessio D’Alì, George Balla, Donggyu Kim, Matt Larson, Dante Luber, Steve Noble, Felipe Rincón, and Cynthia Vinzant for helpful conversations. In the course of this project the second author was supported by Heilbronn Institute for Mathematical Research. The third author received support from the Engineering and Physical Sciences Research Council [grant number EP/X001229/1], as well as the Institute for Advanced Study School of Mathematics and the Fields Institute for Research in Mathematical Sciences. The fourth author received support from the Engineering and Physical Sciences Research Council [grant number EP/X036723/1].
@@ 1.1 Comparison to other work
There has been much subsequent interest in interactions between (type $A$) matroids and Grassmannians, [Ardila2018, Section 4.1] being a survey. When later authors have paid attention to the precise shape of the Grassmann–Plücker equations it has often been because they were generalising matroids, whether by replacing the Grassmannian, as below, or the coefficients. Particularly clear is the discussion in the introduction of [DressWenzel1992], the work introducing valuated matroids, where the coefficients are real numbers representing the images of elements of $\mathbb{K}$ under a nonarchimedean valuation. Valuated matroids found a home in the field of tropical geometry, where they parametrise tropical linear spaces [Speyer2008, Remark 2.1], and the coefficient object is interpreted as the tropical semifield $\mathbb{T}=\mathbb{R}\cup\{\infty\}$. It’s because of this line of work that we have adopted the language of tropicalising equations, even though the solutions to tropical equations that we consider are valued in $\{0,\infty\}$ rather than in all of $\mathbb{T}$.
@@
Valuated matroids can also be defined to be in bijection with regular subdivisions of matroid polytopes into matroid polytopes. We will not use this perspective, but it appears in several of the works below, and its generalisation to strong Coxeter matroids for minuscule representations will be the theme of the forthcoming work [Frankfurt].
@@
In the other classical types $B$, $C$, and $D$, there are fewer minuscule representations than in type $A$. The interesting ones come from the long root in type $B$ and the two roots at the forked end of the Dynkin diagram in type $D$. The Coxeter matroids for both these types were introduced by Bouchet [Bouchet1987] as $\Delta$-matroids, as well as at similar dates by other authors (see [Noble2025, Section 1] for a survey). What a realisation of a $\Delta$-matroid is, and therefore how to coordinatise $\Delta$-matroids, is a hairier question than for matroids; from our perspective this is because of the potential to associate the geometry in any of the types $B$, $C$, or $D$, or even other alternatives, like replacing the bilinear forms in these geometries with sesquilinear ones. Declaring a realisation to be a point of $\mathbb{G}/\mathbb{P}$, as we do in this paper, is the approach of [BorovikGelfandWhite2003], followed e.g. by [Booth2001] introducing oriented $\Delta$-matroids. In the structural literature, realisations are mostly taken to be symmetric or skew-symmetric square matrices, as in [GeelenThesis]. Many authors, following [BouchetDuchamp], further restrict attention to $\mathbb{F}_{2}$, symmetric and skew-symmetric matrices then being no different. From our perspective, these square matrices should be thought of as obtained from a Lagrangian subspace for a skew-symmetric or symmetric quadratic form (such subspaces being what $\mathbb{G}/\mathbb{P}$ parametrises) by taking a matrix spanning the subspace and deleting a maximal identity submatrix.
@@
The first tropical investigations outside type $A$ were made by Rincón [Rincon], who considered even $\Delta$-matroids (type $D$). Here the Wick relations generate the quadric equations of $\mathbb{G}/\mathbb{P}$. [Rincon, Theorem 5.4] is the regular subdivision statement connecting even $\Delta$-matroids to the Wick relations. Our Proposition 3.7 is not a consequence of Rincón’s theorem because he takes the support being a $\Delta$-matroid as an assumption. Rincón also states results about the tropicalisation of $\mathbb{G}/\mathbb{P}$ (i.e. of not just its quadrics), which from the $\Delta$-matroidal point of view is a question of realisability, and about a definition of “isotropic” applying to tropical linear spaces. Balla and Olarte [BallaOlarte] investigate these latter two themes in all symplectic Grassmannians, the case of maximal parabolics in type $C$.
@@ 2.1.1 Coxeter groups, root systems and Weyl groups
Let $\mathcal{V}$}} be an $n$-dimensional Euclidean real vector space with inner product $(\cdot,\cdot)$. For a given hyperplane $H\subset{{\mathcal{V}}}$, we let $s_{H}$ denote the reflection which fixes $H$ pointwise and sends a normal vector of $H$ to its opposite. A Coxeter group $W$ is a finite group generated by a set of reflections in the orthogonal group $O({{\mathcal{V}}})$. A reflecting hyperplane for $W$ is a hyperplane $H$ such that $s_{H}\in W$. The set of reflecting hyperplanes of $W$ forms the Coxeter arrangement of $W$.
For a given $W$ we consider a root system $\Phi\subset{{\mathcal{V}}}$, i.e. a set of vectors closed under the action of $W$, containing one opposite pair of normals to each of the reflecting hyperplanes of $W$. Any choice of linear functional on $\mathcal{V}$}} that’s nonzero on every root separates the root system into a positive part $\Phi^{+}$ and a negative part $\Phi^{-}$. The elements in these parts are called positive roots and negative roots respectively. We let $\Pi\subseteq\Phi^{+}$ be a minimal set with the property that every positive root is a positive linear combination of roots in $\Pi$. The elements of $\Pi$ are called simple roots. We let $S\subseteq W$ consist of the reflections which have a simple root as a normal vector and name these the simple reflections. Then $S$ generates $W$.
@@ 2.2 Ordered (multi)sets and (multi)permutations
For two ordered multisets $A$ and $B$, we let $A\circ B$ denote their concatenation as an ordered multiset. For an ordered multiset $A=\{A_{1},\ldots,A_{m}\}$ we let ${A}^{T}=\{A_{m},\ldots,A_{1}\}$ denote the reversal of $A$. For an ordered set $A$ and a set $B$, we let ${A}\!\setminus\!B$ be the ordered set obtained by deleting the elements of $A\cap B$ from $A$, keeping the relative order of the rest. As an abuse of notation, we use ${A}\!\setminus\!i$ to mean ${A}\!\setminus\!\{i\}$. Additionally, if $i\in A$ we let $A(i)$ denote the index of $i$ in $A$. Note that $A{A}\!\setminus\!\{i\}$ and $A(i)$ are not defined for general ordered multisets since $i$ might appear more than once.
@@
Though this notation suffices throughout, we do distinguish a special family of multipermutations. Given two ordered sets $A$ and $B$, we write $\sigma_{A,B}:=\sigma_{A\circ B}$ for the multipermutation that sorts $A\circ B$ into weakly increasing order, and accordingly $\ell\left(\sigma_{A,B}\right)=\ell\left(A\circ B\right)$ for its length. If we write $\sigma_{A,B}$ for sets $A$ and $B$ to which we have not given an ordering, we mean $A$ and $B$ to be increasingly ordered.
@@ 2.3.1 Coxeter matroids from polytopes
Figure 2: The Coxeter arrangement of the type $B_{3}$ Coxeter group. The edges labelled by $s_{1}$ are solid black, those labelled by $s_{2}$ are dashed red and those labelled by $s_{3}$ are dotted blue.
@@ 3 Characterising strong Coxeter matroids
In this section we give combinatorial characterisations of strong minuscule Coxeter matroids $M$, type by type. At least for the classical types, similar results are already in the literature, but we want characterisations of a special form. We only use conditions on the polytope $P(M)$ that consider whether pairs of antipodal vertices on a given face $Q$ of the ambient polytope $\operatorname{conv}\{W\cdot\omega_{J}\}$ can be present, in particular by requiring that among a chosen set of antipodes in $Q$, not exactly one pair of antipodal vertices can appear in $P(M)$. The reason for preferring these particular conditions is that they can be cast as tropical equations, defined below, that are suitably homogeneous. Then in Section 4 we will see the same tropical equations by tropicalising equations for $\mathbb{G}/\mathbb{P}$.
@@ 3.1.1 Tropical equations
When we call an element $f\in\mathbb{B}[X]$ a tropical equation, we are thinking of it as a condition that may or may not be satisfied by a tuple $p\in\mathbb{B}^{X}$. The tropical equation $f=\bigoplus_{a\in A}X^{\odot a}$ is satisfied by $p$ if the number of $a\in A$ such that $p^{\odot a}=1$ – i.e. such that $p(x)=1$ for all $x$ with $a(x)>0$ – is not exactly 1. If $F\in\mathbb{C}[X]$ is a complex polynomial such that $F(P)=0$ at the point $P\in\mathbb{C}^{X}$, and we let $f=\operatorname{trop}(F)$ and $p\in\mathbb{B}^{X}$ be the support of $P$, i.e. $p(x)=1$ if and only if $P(x)\neq 0$, then the tropical equation $f$ is satisfied by $p$.
@@ 3.2.2 Even $\Delta$-matroids (Type $D$)
The type $D$ strong exchange equations are a subset of the tropical Wick relations [Rincon, Definition 4.2], where one defines $f^{(D)}_{I,J}$ for all $I,J\subseteq[n]$ and not just those of the same parity. Adding in these extra equations allows one to drop the assumption from Proposition 3.7 that $M$ is already an even set system. A special case of [Rincon, Theorem 5.4] states that an arbitrary set system $M$ is an even $\Delta$-matroid if and only if $\nu_{M}$ satisfies the tropical Wick relations.

### v2

@@ 1 Introduction
Many have studied (see Section 1.1) how well this happy state of affairs replicates for other quotients $\mathbb{G}$$\mathbb{P}$} of a reductive group by a parabolic subgroup. In these studies, the first choice to be made is an embedding of $\mathbb{G}$$\mathbb{P}$} in a projective space with a prescribed coordinate basis, which is required to define the tropicalisation. Our objective in this paper is to do so uniformly in $\mathbb{G}$}. Representation theory gives a uniform procedure to define a projective embedding of $\mathbb{G}$$\mathbb{P}$} in a projectivised $\mathbb{G}$}-representation $\operatorname{proj}(V_{\lambda})$, in which it is cut out by quadrics (Theorem 4.2). As for the coordinate basis, for Grassmannians we can concisely express the choice as the basis of homogeneous linear forms in the weight grading for $\mathbb{G}$}; since the weight spaces are one-dimensional, this uniquely specifies our basis up to scalars. The same approach works for $\mathbb{G}$$\mathbb{P}$} when $V_{\lambda}$ is minuscule (we also call $\mathbb{P}$} minuscule in this case), i.e. all the weights of $V_{\lambda}$ are in a single Weyl group orbit. This implies that all weight spaces are one-dimensional. So it is the minuscule case we study here. For convenience in the representation theory, our paper will set $\mathbb{K}$$\mathbb{C}$} (although see Remark 4.1).
On the other side of the connection, Coxeter matroids [BorovikGelfandWhite2003] are to $\mathbb{G}$$\mathbb{P}$} as matroids are to the Grassmannian. A Coxeter matroid is any subpolytope of the orbit polytope associated to $\mathbb{P}$} of the Weyl group $W({\mathbb{G}})$ whose edges are parallel to roots of $\mathbb{G}$} (see Section 2.3 for a careful definition). With this, we have every piece but one of our main theorem:
@@
The one word yet to be defined is “strong”. We say that a Coxeter matroid $P$ is strong if for any two vertices $v,w$ of $P$, either $\operatorname{conv}\{v,w\}$ is parallel to a root or it is the diagonal of a parallelogram all four of whose vertices are vertices of $P$ and with an edge parallel to a root. (Such Coxeter matroids are also said to satisfy the strong exchange property. For more details see Section 2.3 below.) The condition of being strong will emerge from the structure of tropicalised quadrics homogeneous in the weight grading. In particular, if $P$ is realizable in the sense that there is a point $x\in{\mathbb{G}}/{\mathbb{P}}$ so that a vertex $v$ of the orbit polytope is present in $P$ if and only if the weight-$v$ coordinate of $x$ is nonzero, then $P$ satisfies every tropical equation of $\mathbb{G}$$\mathbb{P}$}, so $P$ is strong. All (ordinary) matroids are strong, but not, for example, all $\Delta$-matroids (Section 2.5). For Coxeter matroids that are not strong, we are unable to suggest any characterisation by equations of tropical hypersurfaces, whatever their degree.
Borovik, Gelfand and White recognised the importance of minuscule representations in Open Problem 6.16.1 of [BorovikGelfandWhite2003]. Our theorem can be read as morally confirming the “if” direction of their expectation in that open problem, strengthened by the connection to equations of $\mathbb{G}$$\mathbb{P}$}.
@@ Acknowledgements
The authors thank Chris Eur for his part in the initial impetus for this project and for pointing us to [Kumar2002], Matt Larson and Igor Makhlin for comments on the isomorphism between type $B_{n}$ and $D_{n+1}$ spinor Grassmannians, Mauricio Velasco for pointers to the history of the Wick relations in the literature, and Alessio D’Alì, George Balla, Donggyu Kim, Dante Luber, Steve Noble, Felipe Rincón, and Cynthia Vinzant for helpful conversations. In the course of this project the second author was supported by Heilbronn Institute for Mathematical Research. The third author received support from the Engineering and Physical Sciences Research Council [grant number EP/X001229/1], as well as the Institute for Advanced Study School of Mathematics and the Fields Institute for Research in Mathematical Sciences. The fourth author received support from the Engineering and Physical Sciences Research Council [grant number EP/X036723/1].
@@ 1.1 Comparison to other work
There has been much subsequent interest in interactions between (type $A$) matroids and Grassmannians, [Ardila2018, Section 4.1] being a survey. When later authors have paid attention to the precise shape of the Grassmann–Plücker equations it has often been because they were generalising matroids, whether by replacing the Grassmannian, as below, or the coefficients. Particularly clear is the discussion in the introduction of [DressWenzel1992], the work introducing valuated matroids, where the coefficients are real numbers representing the images of elements of $\mathbb{K}$} under a nonarchimedean valuation. Valuated matroids found a home in the field of tropical geometry, where they parametrise tropical linear spaces [Speyer2008, Remark 2.1], and the coefficient object is interpreted as the tropical semifield $\mathbb{T}$$\mathbb{R}$$\cup\{\infty\}$. It’s because of this line of work that we have adopted the language of tropicalising equations, even though the solutions to tropical equations that we consider are valued in $\{0,\infty\}$ rather than in all of $\mathbb{T}$}.
@@
Valuated matroids can also be defined to be in bijection with regular subdivisions of matroid polytopes into matroid polytopes. We will not use this perspective, but it appears in several of the works below, and its generalisation to strong Coxeter matroids for minuscule representations is the theme of [Frankfurt].
@@
In the other classical types $B$, $C$, and $D$, there are fewer minuscule representations than in type $A$. The interesting ones come from the long root in type $B$ and the two roots at the forked end of the Dynkin diagram in type $D$. The Coxeter matroids for both these types were introduced by Bouchet [Bouchet1987] as $\Delta$-matroids, as well as at similar dates by other authors (see [Noble2025, Section 1] for a survey). What a realisation of a $\Delta$-matroid is, and therefore how to coordinatise $\Delta$-matroids, is a hairier question than for matroids; from our perspective this is because of the potential to associate the geometry in any of the types $B$, $C$, or $D$, or even other alternatives, like replacing the bilinear forms in these geometries with sesquilinear ones. Declaring a realisation to be a point of $\mathbb{G}$$\mathbb{P}$}, as we do in this paper, is the approach of [BorovikGelfandWhite2003], followed e.g. by [Booth2001] introducing oriented $\Delta$-matroids. In the structural literature, realisations are mostly taken to be symmetric or skew-symmetric square matrices, as in [GeelenThesis]. Many authors, following [BouchetDuchamp], further restrict attention to ${\mathbb{F}}_{2}$, symmetric and skew-symmetric matrices then being no different, at least off the diagonal. From our perspective, these square matrices should be thought of as obtained from a Lagrangian subspace for a skew-symmetric or symmetric quadratic form (such subspaces being what $\mathbb{G}$$\mathbb{P}$} parametrises) by taking a matrix spanning the subspace and deleting a maximal identity submatrix.
@@
The first tropical investigations outside type $A$ were made by Rincón [Rincon], who considered even $\Delta$-matroids (type $D$). Here the Wick relations generate the quadric equations of $\mathbb{G}$$\mathbb{P}$}. The history of the Wick relations in the literature is complicated; see Remark 4.12 for our attempts to trace it. In [Rincon] the tropical Wick relations appear as the definition of valuated even $\Delta$-matroid, without a direct comparison to the unvaluated case. [Rincon, Theorem 5.4] is the regular subdivision statement connecting even $\Delta$-matroids to the Wick relations, but our Proposition 3.7 is not a consequence of Rincón’s theorem because he takes the support being a $\Delta$-matroid as an assumption. Rincón also states results about the tropicalisation of $\mathbb{G}$$\mathbb{P}$} (i.e. of not just its quadrics), which from the $\Delta$-matroidal point of view is a question of realisability, and about a definition of “isotropic” applying to tropical linear spaces. Balla and Olarte [BallaOlarte] investigate these latter two themes in all symplectic Grassmannians, the case of maximal parabolics in type $C$.
@@ 2.1.1 Coxeter groups, root systems and Weyl groups
Let $\mathcal{V}$}} be an $n$-dimensional Euclidean real vector space with inner product $(\cdot,\cdot)$. For a given hyperplane $H\subset{{\mathcal{V}}}$, we let $s_{H}$ denote the reflection which fixes $H$ pointwise and sends a normal vector of $H$ to its opposite. A finite Coxeter group $W$ is a finite group generated by a set of reflections in the orthogonal group $O({{\mathcal{V}}})$. A reflecting hyperplane for $W$ is a hyperplane $H$ such that $s_{H}\in W$. The set of reflecting hyperplanes of $W$ forms the Coxeter arrangement of $W$.
For a given $W$ we consider a root system $\Phi\subset{{\mathcal{V}}}$, i.e. a set of vectors closed under the action of $W$, containing one opposite pair of normals to each of the reflecting hyperplanes of $W$. Any choice of linear functional on $\mathcal{V}$}} that’s nonzero on every root separates the root system into a positive part $\Phi^{+}$ and a negative part $\Phi^{-}$. The elements in these parts are called positive roots and negative roots respectively. We let $\Pi\subseteq\Phi^{+}$ be a minimal set with the property that every positive root is a positive linear combination of roots in $\Pi$. The elements of $\Pi$ are called simple roots. We let $S\subseteq W$ consist of the reflections which have a simple root as a normal vector to its reflecting hyperplane and name these the simple reflections. Then $S$ generates $W$.
@@ 2.2 Ordered (multi)sets and (multi)permutations
For two ordered multisets $A$ and $B$, we let $A\circ B$ denote their concatenation as an ordered multiset. For an ordered multiset $A=\{A_{1},\ldots,A_{m}\}$ we let ${A}^{T}=\{A_{m},\ldots,A_{1}\}$ denote the reversal of $A$. For an ordered set $A$ and a set $B$, we let ${A}\!\setminus\!B$ be the ordered set obtained by deleting the elements of $A\cap B$ from $A$, keeping the relative order of the rest. As an abuse of notation, we use ${A}\!\setminus\!i$ to mean ${A}\!\setminus\!\{i\}$. Additionally, if $i\in A$ we let $A(i)$ denote the index of $i$ in $A$. Note that ${A}\!\setminus\!\{i\}$ and $A(i)$ are not (well-)defined for general ordered multisets since $i$ might appear more than once.
@@
Though this notation suffices throughout, we do distinguish a special family of multipermutations. Given two ordered sets $A$ and $B$, for readability we write $\sigma_{A,B}:=\sigma_{A\circ B}$ for the multipermutation that sorts $A\circ B$ into weakly increasing order, and accordingly $\ell\left(\sigma_{A,B}\right)=\ell\left(A\circ B\right)$ for its length. If we write $\sigma_{A,B}$ for sets $A$ and $B$ to which we have not given an ordering, we mean $A$ and $B$ to be increasingly ordered.
@@ 2.3.1 Coxeter matroids from polytopes
Figure 2: The Coxeter arrangement of the type $B_{3}$ Coxeter group. The edges labelled by $s_{1}$ are solid black, those labelled by $s_{2}$ are dashed red and those labelled by $s_{3}$ are dotted blue. The labels represent right multiplication.
@@ 3 Characterising strong Coxeter matroids
In this section we give combinatorial characterisations of strong minuscule Coxeter matroids $M$, type by type. At least for the classical types, similar results are already in the literature, but we want characterisations of a special form. We only use conditions on the polytope $P(M)$ that consider whether pairs of antipodal vertices on a given face $Q$ of the ambient polytope $\operatorname{conv}\{W\cdot\omega_{J}\}$ can be present, in particular by requiring that among a chosen set of antipodes in $Q$, not exactly one pair of antipodal vertices can appear in $P(M)$. The reason for preferring these particular conditions is that they can be cast as tropical equations, defined below, that are suitably homogeneous. Then in Section 4 we will see the same tropical equations by tropicalising equations for $\mathbb{G}$$\mathbb{P}$}.
@@ 3.1.1 Tropical equations
When we call an element $f\in{\mathbb{B}}[X]$ a tropical equation, we are thinking of it as a condition that may or may not be satisfied by a tuple $p\in{\mathbb{B}}^{X}$. The tropical equation $f=\bigoplus_{a\in A}X^{\odot a}$ is satisfied by $p$ if the number of $a\in A$ such that $p^{\odot a}=1$ – i.e. such that $p(x)=1$ for all $x$ with $a(x)>0$ – is not exactly 1. If $F\in{\mathbb{C}}[X]$ is a complex polynomial such that $F(P)=0$ at the point $P\in{\mathbb{C}}^{X}$ and we let $f=\operatorname{trop}(F)$ and $p\in{\mathbb{B}}^{X}$ be the support of $P$, i.e. $p(x)=1$ if and only if $P(x)\neq 0$, then the tropical equation $f$ is satisfied by $p$.
@@ 3.2.2 Even $\Delta$-matroids (Type $D$)
The type $D$ strong exchange equations are a subset of the tropical Wick relations [Rincon, Definition 4.2], where one defines $f^{(D)}_{I,J}$ for all $I,J\subseteq[n]$ and not just those of the same parity. Adding in these extra equations allows one to drop the assumption from Proposition 3.7 that $M$ is already an even set system.
@@
Remark 3.10.
One can construct a map from the vertices of the $n$-cube $Q_{n}$ to the even (or odd) demicube ${\mathcal{D}}_{n+1}^{+}$ (${\mathcal{D}}_{n+1}^{-}$): the map to the even demicube is
$I\mapsto\begin{cases}I&|I|\text{ even,}\\ I\cup\{n+1\}&|I|\text{ odd}\end{cases}$
with the odd version defined similarly. [Murota2021, Theorem 3.2] shows that a set system is a type $B_{n}$ strong Coxeter matroid if and only if its image under this map is a type $D_{n+1}$ (strong) Coxeter matroid. This can also be deduced from the strong exchange equations: under this map, the type $B$ strong exchange equations (14) and (15) map to the type $D$ strong exchange equations (17).